# Emergent broadband polarization entanglement from electronic and phononic Stokes–anti-Stokes indistinguishability

### v2

@@ Abstract
Recently [PRA 108, L051501 (2023)], it has been shown that in a centrosymmetric cubic system, two-photons from a broadband intense laser field can be converted into a pair of Stokes and anti-Stokes (SaS) entangled photons. While the previous work was based on symmetry arguments, here we present a fully quantum theory for the SaS scattering that properly explains, quantitatively describes, and provides a means to predict its spectral and polarization properties (for diamond). We also explore the possibilities offered by such system, designing an entanglement map based on changes in the light-matter system. In particular, we show how the broadband polarization entanglement, that emerges from the interference between electronic and phononic degrees of freedom in the SaS scattering, depends on parameters such as Stokes-anti-Stokes Raman shift, scattering geometry and laser bandwidth, opening the avenue of exploration of such phenomenon in information processing.
Photons produced in Raman scattering have their frequency shifted in relation to the incoming light ($\omega_{L}$), being called Stokes if their frequency is red-shifted ($\omega_{S}$) and anti-Stokes if they are blue-shifted ($\omega_{aS}$). The two of them can be produced as part of a higher order process, which results in their correlation ($\hbar\omega_{L}+\hbar\omega_{L}^{\prime}=\hbar\omega_{S}+\hbar\omega_{aS}$), with the coupled quanta being known as the SaS photon pair [1, 2, 3]. The operation of converting incoming photon(s) into different outgoing photon(s) in a medium is dictated by its susceptibility $\chi$ [4, 5], and the $\chi^{(3)}$ enables two photons from a broadband intense laser field to interact with the medium to produce the correlated SaS photon pair. In centrosymmetric materials, the second-order susceptibility is null and the non-linear response is usually dominated by the third-order tensor [6]. While the non-classical correlations between SaS photons have been widely established [7, 8, 9, 10, 11], the presence of polarization entanglement in such process has been demonstrated only very recently, in diamond [12].
@@
The correlated SaS scattered intensity can be obtained by the electric polarization in the material [4, 5]
@@
[…]
@@
×$\times\left[A^{E}_{ijkl}+A^{R}_{ijkl}\frac{\gamma}{\omega_{ph}-\omega+i\gamma/2}\right]d\omega,$ (2)
where the total SaS susceptibility (in brackets) is composed of a constant electronic and a resonant phononic component [5, 14, 4].
The diamond structure belongs to the $O_{h}$ point group [13, 14]. In this case, $\chi^{(3)}_{ijkl}$ has only four independent terms [6], which are $\chi^{(3)}_{xxxx},\chi^{(3)}_{xxyy},\chi^{(3)}_{xyxy},\chi^{(3)}_{xyyx}$ (assuming light propagating in the $z$ direction and that the $x$ and $y$ directions coincide with the crystallographic axes). The frequency dispersion of the electronic contribution e-SaS can be neglected when it is very far from any electronic resonance, so that $\chi^{(3)E}_{ijkl}(-\omega_{aS},\omega_{L},\omega_{L}^{\prime},-\omega_{S})=A^{E}_{ijkl}\delta(\omega_{L}+\omega_{L}^{\prime}-\omega_{S}-\omega_{aS})$ becomes a complex constant tensor fulfilling energy conservation. Ideally, the ratio between the e-SaS components in centrosymmetric materials is $\chi^{(3)E}_{xxxx}\approx 3\chi^{(3)E}_{xyyx}\approx 3\chi^{(3)E}_{xxyy}\approx 3\chi^{(3)E}_{xyxy}$ [4]. However, experimental conditions are rarely ideal and we will only take these values as a reference.
The p-SaS susceptibility tensor, on the other hand, is known to be $\chi^{(3)R}_{ijkl}\propto\sum_{\sigma}(\alpha^{R}_{ij,\sigma}\alpha^{R}_{kl,\sigma}+\alpha^{R}_{ik,\sigma}\alpha^{R}_{jl,\sigma})$ [13], where $\alpha^{R}_{ij,\sigma}$ is the polarizability Raman tensor that describes the scattering of an incident electric field polarized at $i$ into a scattered mode polarized at $j$, via a phonon $\sigma$. The Raman active vibrational mode of diamond belongs to the $T_{2g}$ irreducible representation of the $O_{h}$ point group. Due to the relatively small probability of formation of SaS pairs, the spectral features of the phononic contribution to the susceptibility can be calculated from perturbation theory [4], and $A^{R}_{ijkl}$ in Eq. (2) is the tensorial part of $\chi^{(3)R}_{ijkl}$. The frequency dependence is composed of a Lorentzian-shaped amplitude of the Stokes scattering phonon frequency, $\omega=\omega_{L}-\omega_{S}$, around the resonance $\omega_{ph}$ with width $\gamma/2$ ($\gamma$ is the phonon decay rate, inversely proportional to its lifetime). This susceptibility is the same as when the scattered fields are classical [5, 4], being an extrapolation of the stimulated Raman scattering to the regime where the Stokes stimulation is turned off.
@@
The SaS scattering amplitude is widened by the laser bandwidth, which adds an energy uncertainty to the scattered photons in both electronic and phononic processes. The factor $C$ in Eq. (2) contains the efficiency of the SaS scattering in polarizations $l$ (Stokes) and $i$ (anti-Stokes), and depends on the scattered field frequencies and the scattering angles. However, we can consider $C$ a constant in our experiments because we only collect pairs in forward scattering and because the laser bandwidth and phonon spectrum, taken care separately in Eq. (2), dominate any other frequency dependence. Assuming a Gaussian amplitude spectrum for the laser, $G(\omega_{L})=(\pi W^{2})^{-1/4}e^{-(\omega_{L}-\omega_{c})^{2}/2W^{2}}$, centered around the angular frequency $\omega_{c}$ with width $W$ (the laser power spectrum FWHM is $2\sqrt{\text{ln}2}W/(2\pi)$), the integral (2) can be solved analytically.
@@
With the probability amplitude (2) we can construct the SaS two-photon state generated by the scattering process. We assume the laser polarization to be vertical (V) with respect to the laboratory, denoted by $\hat{\mathbf{e}}_{V}$. If the crystallographic axis $x$ is also along the vertical direction, then $\hat{\mathbf{e}}_{V}=\hat{\mathbf{x}}$, and the susceptibility $\chi^{(3)}_{ixxl}$ will govern the photon pair production. In general, however, if the crystallographic axis $x$ and the laser polarization have an angle $\theta$ between them, such that $\hat{\mathbf{e}}_{V}=\cos(\theta)\hat{\mathbf{x}}+\sin(\theta)\hat{\mathbf{y}}$, there will be a coherent sum of certain components of the susceptibility tensor defining the scattered photons polarization. By writing the laser field components as $\mathcal{E}_{0x}=\mathcal{E}_{0}\cos\theta$ and $\mathcal{E}_{0y}=\mathcal{E}_{0}\sin\theta$, the scattered (non-normalized) state can be written as
@@
$+[Y^{E}_{VH}(\theta)f^{E}+Y^{R}_{VH}(\theta)f^{R}](|VH\rangle+|HV\rangle),$ (3)
where $f^{\eta}(\omega_{S},\omega_{aS})$ carries the spectral dependence of the scattering for $\eta=\{E,R\}$, being $f^{E}\equiv e^{-\bar{\omega}^{2}/W^{2}}$ and
with $\bar{\omega}\equiv(\omega_{S}+\omega_{aS})/2-\omega_{c}$ and $\Omega\equiv(\omega_{aS}-\omega_{S})/2-\omega_{ph}$. Note that Eq. (4) comes from the integral in Eq. (2), and the complex denominator in the Raman term yields the complex argument in the error function, associated with the phonon decay. Also note that in $f^{R}$ the $e^{-\bar{\omega}^{2}/W^{2}}$ factor implied that detections at $\omega_{S}+\omega_{aS}=2\omega_{c}$ are favored, while the dependence with $\Omega$ means that what dictates the spectrum is the SaS combined frequency $(\omega_{aS}-\omega_{S})/2$ around $\omega_{ph}$. The functions $Y^{\eta}_{li}(\theta)$ depend on the angle $\theta$ and the $A^{\eta}_{ijkl}$ tensor components,
@@
×$\times\left(A^{\eta}_{xxxx}-A^{\eta}_{xyyx}-A^{\eta}_{xxyy}-A^{\eta}_{xyxy}\right),$
where $\mathcal{S}\equiv\sin\theta$ and $\mathcal{C}\equiv\cos\theta$. Equation (3) means that the angle $\theta$ drives the balance between the contribution of the e-SaS spectrum in $f^{E}$ and the p-SaS spectrum in $f^{R}$, and this balance will be different for each pair of scattered photons polarization $VV$, $HH$ and $VH$ (equal to $HV$), depending on how the tensor components are combined. If a different, possibly more complicated, laser spectrum were used, a new spectral function $f^{R}$ would result, but the state of Eq. (3) would still have the same form.
@@
The experimental setup utilized is similar to the one described in [12], with the addition of a monochromator to filter the Stokes signal [18]. The polarization of the laser is vertical (V) with respect to the laboratory and, for each monochromator position (each selected $\delta\omega=\omega_{c}-\omega_{S}$), one histogram of temporal difference is obtained in $300$ seconds of acquisition. This scanning procedure is done for two orientations of the sample ($\theta=0^{\circ}$ and $45^{\circ}$), and for each sample orientation the spectrum is obtained with the polarization of the photons incident on the avalanche photodiode (APD) detectors being selected in two ways—a spectrum with Stokes and anti-Stokes photons with vertical polarization (VV) and a spectrum with both photons with horizontal polarization (HH). The sample utilized was a diamond grown by a CVD process (Type IIac, 100-oriented, from Almax) positioned so that the laser propagates in the (001) direction of the crystal.
@@
The intensity spectra of correlated SaS pairs for VV (blue dots) and HH (red dots) polarization at $\theta=0^{\circ}$ and $\theta=45^{\circ}$ are displayed in Fig. 1 (a) and (b), respectively. This experiment was specifically designed by combining the single-wavelength polarization analysis presented in [12] with the unpolarized spectral analysis provided in [15], to prove the strength of our theoretical approach. According to group theory analysis, VV(0) configuration does not exhibit a Raman contribution ($A^{R}_{xxxx}=0$ in Eq. (5)) and only involves non-resonant electronic transitions. For this reason, the blue curve shown in Fig. 1 (a) corresponding to this configuration is flat. Additionally, Fig. 1 (a) shows a light blue solid line, obtained by averaging the intensity values $I_{SaS}^{corr}$ for this experimental configuration, $\langle I_{SaS}(VV(0^{\circ})\rangle=(27.5\pm 3.5)\times 10^{3}$. Since $Y^{E}_{VV}(0^{\circ})=C\mathcal{E}_{0}^{2}A^{E}_{xxxx}$ and in our model it is proportional to $\langle I_{SaS}(VV(0^{\circ})\rangle^{1/2}$, we have a reference value for $A^{E}_{xxxx}$.
@@
By looking at the other relevant scattering geometries, we obtain the other tensorial components of Eq. (5) in relation to $A^{E}_{xxxx}$. The HH(45) configuration also exhibits only e-SaS, as seen in the red data in Fig 1 (b). The mean and standard deviation of the counts gives $(1.60\pm 0.28)\times 10^{3}$. The HH(0) data in Fig. 1 (a) (red points) contains p-SaS with the characteristic resonance signature, which comes from $A^{R}_{xyyx}\neq 0$ in Eq. (5b), contributing with a Raman scattering in the $|HH\rangle$ polarization. Using $Y^{E}_{HH}(0^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(0.68-i0.12)$ and $Y^{R}_{HH}(0^{\circ})=51450$ fits the data with the shown theoretical curve in Fig. 1 (a). The VV(45) data is the last configuration left, and it is shown in Fig. 1 (b) in blue. In order to fit it, we fix the Raman factor to be the same as in the HH(0) configuration, $Y^{R}_{VV}(45^{\circ})=51450$, and use $Y^{E}_{VV}(45^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(1.61-i0.55)$. With the above values and working with Eq. (5) we obtain the electronic and Raman $A^{\eta}_{ijkl}$ components, summarized in Table 1.
@@
Table 1: List of measured electronic and Raman tensor components ratios $A^{\eta}_{ijkl}/A^{E}_{xxxx}$ [18].
@@
$(A^{E}_{xxyy}+A^{E}_{xyxy})/A^{E}_{xxxx}$ $(0.89\pm 0.0.10)+i(-0.07\pm 0.01)$
$A^{R}_{xxxx}/A^{E}_{xxxx}$ 00
@@
For completeness the second-order correlation function $g^{(2)}(0)$ is evaluated with the ratio $[I^{corr}_{SaS}(\Delta\tau=0)+\bar{I}_{SaS}(\Delta\tau\neq 0)]/\bar{I}_{SaS}(\Delta\tau\neq 0)$, where $I^{corr}_{SaS}(\Delta\tau=0)$ are the measurements in Fig. 1 (a) and (b), and $\bar{I}_{SaS}(\Delta\tau\neq 0)$ accounts for the uncorrelated pair production. A plot of $g^{(2)}(0)$ is shown in Fig. 1(c,d), where it is evident that near the resonance the values are the lowest, accounting for the high number of uncorrelated SaS photon pairs produced. The large error bars in $g^{(2)}(0)$ reflect the errors in the low accidental counts (see [18]), and the theory agrees well with the data in spite of the error size. Furthermore, the curves for HH(0) and VV(45) show a drop in $g^{(2)}(0)$, being higher below and lower above the resonance region, while the VV($0^{\circ}$) and HH($45^{\circ}$) keep roughly the same value. The asymmetry in HH(0) and VV(45) is associated with the uncorrelated counts being symmetrical with respect to the resonance peak, and thus the asymmetry in the correlated counts for these configurations, which contain a p-SaS contribution, is transferred to the $g^{(2)}(0)$ curves. This asymmetry is a result of the constructive ($\delta\omega<\omega_{ph}$) versus destructive ($\delta\omega>\omega_{ph}$) interference between the e-SaS and the p-SaS, which explains the Cooper-pair-like behavior of $I_{SaS}$ [10, 15]. On the other hand, the correlated counts in the VV(0) and HH(45) configurations are a purely e-SaS contribution, and thus are a flat curve, leading to an overall symmetric $g^{(2)}(0)$.

### v3

@@ Abstract
Recently [PRA 108, L051501 (2023)], it was shown that in a centrosymmetric cubic system, two photons from a broadband intense laser field can be converted into a pair of Stokes and anti-Stokes (SaS) entangled photons. While the previous work was based on symmetry arguments, here we present a fully quantum theory for the SaS scattering that properly explains, quantitatively describes, and provides a means to predict its spectral and polarization properties (for diamond). We also explore the possibilities offered by such system, designing an entanglement map based on changes in the light-matter system. In particular, we show how the broadband polarization entanglement, that emerges from the interference between electronic and phononic degrees of freedom in the SaS scattering, depends on parameters such as Stokes-anti-Stokes Raman shift, scattering geometry and laser bandwidth, opening the avenue of exploration of such phenomenon in information processing.
I Introduction
Photons produced in Raman scattering have their frequency shifted in relation to the incoming light ($\omega_{L}$), being called Stokes if their frequency is red-shifted ($\omega_{S}$) and anti-Stokes if they are blue-shifted ($\omega_{aS}$). The two of them can be produced as part of a higher order process, which results in their correlation ($\hbar\omega_{L}+\hbar\omega_{L}^{\prime}=\hbar\omega_{S}+\hbar\omega_{aS}$), with the coupled quanta being known as the Stokes–anti-Stokes (SaS) photon pair [1, 2, 3]. The operation of converting incoming photon(s) into different outgoing photon(s) in a medium is dictated by its susceptibility $\chi$ [4, 5], and the $\chi^{(3)}$ enables two photons from a broadband intense laser field to interact with the medium to produce the correlated SaS photon pair. In centrosymmetric materials, the second-order susceptibility is null and the non-linear response is usually dominated by the third-order tensor [6]. While the non-classical correlations between SaS photons have been widely established [7, 8, 9, 10, 11], the presence of polarization entanglement in such process has been demonstrated only very recently, in diamond [12].
@@ I Introduction
Our article is organized as follows. In Sec. II we provide the theory to support our experiments in which the two-photon quantum state arising in the Stokes–anti-Stokes scattering is derived. Section III describes the experimental methods and their results, including the measurement of the Raman tensor components via spectral measurements for different photon polarization pairs and crystal angles, which is obtained with our theory. In Sec. IV, the polarization entanglement of the SaS photon pairs is discussed, and we construct entanglement maps showing where to find maximum entanglement within the experimental parameters. We conclude the article in Sec. V.
II Two-photon SaS quantum state
The correlated SaS scattered intensity can be obtained by the third-order electric polarization in the material [4, 5]
@@ II Two-photon SaS quantum state
where $\mathcal{E}$ and $E$ represent the laser and scattered modes, respectively, and indexes $\{i,j,k,l\}$ represent polarization directions. The laser modes are occupied by an intense classical field, which is unaffected by the scattering of a few photons, while the Stokes and anti-Stokes modes start in the vacuum. The incident laser spectral amplitude is written as $\mathcal{E}_{j}(\omega_{L})=\mathcal{E}_{0j}G(\omega_{L})$, with $G(\omega_{L})$ describing the laser spectral distribution with $\int_{0}^{\infty}|G(\omega_{L})|^{2}d\omega_{L}=1$.
The total SaS susceptibility is composed of an electronic (e-SaS) and a phononic (p-SaS) component [5, 14, 4], $\chi^{(3)}_{ijkl}=\chi^{(3)E}_{ijkl}+\chi^{(3)R}_{ijkl}$, with $E$ for electronic and $R$ for Raman phononic. The diamond structure belongs to the $O_{h}$ point group [13, 14], in which case $\chi^{(3)}_{ijkl}$ has only four independent terms [6], which are $\chi^{(3)}_{xxxx},\chi^{(3)}_{xxyy},\chi^{(3)}_{xyxy},\chi^{(3)}_{xyyx}$ (assuming light propagating in the $z$ direction and that the $x$ and $y$ directions coincide with the crystallographic axes). The frequency dispersion of the electronic contribution e-SaS can be neglected when it is very far from any electronic resonance, so that $\chi^{(3)E}_{ijkl}(-\omega_{aS},\omega_{L},\omega_{L}^{\prime},-\omega_{S})=A^{E}_{ijkl}\delta(\omega_{L}+\omega_{L}^{\prime}-\omega_{S}-\omega_{aS})$ becomes a complex constant tensor fulfilling energy conservation. Ideally, the ratio between the e-SaS components in centrosymmetric materials is $\chi^{(3)E}_{xxxx}\approx 3\chi^{(3)E}_{xyyx}\approx 3\chi^{(3)E}_{xxyy}\approx 3\chi^{(3)E}_{xyxy}$ [4]. However, experimental conditions are rarely ideal and we will only take these values as a reference.
The p-SaS susceptibility tensor, on the other hand, is known to be $\chi^{(3)R}_{ijkl}\propto\sum_{\sigma}(\alpha^{R}_{ij,\sigma}\alpha^{R}_{kl,\sigma}+\alpha^{R}_{ik,\sigma}\alpha^{R}_{jl,\sigma})$ [13], where $\alpha^{R}_{ij,\sigma}$ is the polarizability Raman tensor that describes the scattering of an incident electric field polarized at $i$ into a scattered mode polarized at $j$, via a phonon $\sigma$. The Raman active vibrational mode of diamond belongs to the $T_{2g}$ irreducible representation of the $O_{h}$ point group. Due to the relatively small probability of formation of SaS pairs, the spectral features of the phononic contribution to the susceptibility can be calculated from perturbation theory [4] as being
$\chi^{(3)R}_{ijkl}(-\omega_{aS},\omega_{L},\omega_{L}^{\prime},-\omega_{S})=A^{R}_{ijkl}\frac{\gamma}{\omega_{ph}-\omega+i\gamma/2}\delta(\omega_{L}+\omega_{L}^{\prime}-\omega_{S}-\omega_{aS}).$ (2)
The tensorial part of the susceptibility is $A^{R}_{ijkl}$, and he frequency dependence is composed of a Lorentzian-shaped amplitude of the Stokes scattering phonon frequency, $\omega=\omega_{L}-\omega_{S}$, around the resonance $\omega_{ph}$ with width $\gamma/2$ ($\gamma$ is the phonon decay rate, inversely proportional to its lifetime). This susceptibility is the same as when the scattered fields are classical [5, 4], being an extrapolation of the stimulated Raman scattering to the regime where the Stokes stimulation is turned off.
The probability amplitude that describes the SaS pair scattered at frequencies $\omega_{S}$ and $\omega_{aS}$ in a polarization pure state $|\psi\rangle$ is obtained by calculating $\langle 0|E_{l}(\omega_{S})E_{i}(\omega_{aS})|\psi\rangle+\langle 0|E_{i}(\omega_{aS})E_{l}(\omega_{S})|\psi\rangle$ [16], and since the electric polarization frequency component yields a scattered field in that same frequency and polarization, the two-photon scattering amplitude probability at the SaS frequencies reads
@@
$\times\left[A^{E}_{ijkl}+A^{R}_{ijkl}\frac{\gamma}{\omega_{ph}-\omega+i\gamma/2}\right]d\omega.$ (3)
@@
The SaS scattering amplitude is widened by the laser bandwidth, which adds an energy uncertainty to the scattered photons in both electronic and phononic processes. The factor $C$ in Eq. (II) contains the efficiency of the SaS scattering in polarizations $l$ (Stokes) and $i$ (anti-Stokes), and depends on the scattered field frequencies and the scattering angles. However, we can consider $C$ a constant in our experiments because we only collect pairs in forward scattering and because the laser bandwidth and phonon spectrum, taken separately in Eq. (II), dominate any other frequency dependence. Assuming a Gaussian amplitude spectrum for the laser, $G(\omega_{L})=(\pi W^{2})^{-1/4}e^{-(\omega_{L}-\omega_{c})^{2}/2W^{2}}$, centered around the angular frequency $\omega_{c}$ with width $W$ (the laser power spectrum FWHM is $2\sqrt{\text{ln}2}W/(2\pi)$), the integral (II) can be solved analytically. By separating the e-SaS and p-SaS parts, $f^{E}\equiv e^{-\bar{\omega}^{2}/W^{2}}$ and
with $\bar{\omega}\equiv(\omega_{S}+\omega_{aS})/2-\omega_{c}$ and $\Omega\equiv(\omega_{aS}-\omega_{S})/2-\omega_{ph}$, then $\Psi_{li}=C\mathcal{E}_{0j}\mathcal{E}_{0k}(A^{E}_{ijkl}f^{E}+A^{R}_{ijkl}f^{R})$. Note that Eq. (4) comes from the integral in Eq. (II), and the complex denominator in the Raman term yields the complex argument in the error function, associated with the phonon decay. Also note that in $f^{R}$ the $e^{-\bar{\omega}^{2}/W^{2}}$ factor implies that detections at $\omega_{S}+\omega_{aS}=2\omega_{c}$ are favored, while the dependence with $\Omega$ means that what dictates the spectrum is the SaS combined frequency $(\omega_{aS}-\omega_{S})/2$ around $\omega_{ph}$.
@@
With the probability amplitude (II) we can construct the SaS two-photon state generated by the scattering process. We assume the laser polarization to be vertical (V) with respect to the laboratory, denoted by $\hat{\mathbf{e}}_{V}$. If the crystallographic axis $x$ is also along the vertical direction, then $\hat{\mathbf{e}}_{V}=\hat{\mathbf{x}}$, and the susceptibility $\chi^{(3)}_{ixxl}$ will govern the photon pair production. In general, however, if the crystallographic axis $x$ and the laser polarization have an angle $\theta$ between them, such that $\hat{\mathbf{e}}_{V}=\cos(\theta)\hat{\mathbf{x}}+\sin(\theta)\hat{\mathbf{y}}$, there will be a coherent sum of certain components of the susceptibility tensor defining the scattered photons polarization. By writing the laser field components as $\mathcal{E}_{0x}=\mathcal{E}_{0}\cos\theta$ and $\mathcal{E}_{0y}=\mathcal{E}_{0}\sin\theta$, the scattered (non-normalized) state can be written as
@@
$+[Y^{E}_{VH}(\theta)f^{E}+Y^{R}_{VH}(\theta)f^{R}](|VH\rangle+|HV\rangle).$ (5)
The functions $Y^{\eta}_{li}(\theta)$ depend on the angle $\theta$ and the $A^{\eta}_{ijkl}$ tensor components,
@@
$\times\left(A^{\eta}_{xxxx}-A^{\eta}_{xyyx}-A^{\eta}_{xxyy}-A^{\eta}_{xyxy}\right),$
where $\mathcal{S}\equiv\sin\theta$ and $\mathcal{C}\equiv\cos\theta$. Equation (5) means that the angle $\theta$ drives the balance between the contribution of the e-SaS spectrum in $f^{E}$ and the p-SaS spectrum in $f^{R}$, and this balance will be different for each pair of scattered photons polarization $VV$, HHHH and $VH$ (equal to $HV$), depending on how the tensor components are combined. If a different, possibly more complicated, laser spectrum were used, a new spectral function $f^{R}$ would result, but the state of Eq. (5) would still have the same form.
@@
III Experimental characterization of the centrosymmetric SaS scattering
III.1 Setup and methods
Figure 1: Depiction of experimental configuration. (M) mirror, (HWP) half-wave plate, (PBS) polarization beam splitter, (FI) faraday isolator, (BP1, BP2) interference bandpass filter, (Ob1) objective for laser focusing on the sample (20x, N.A. 0.50, W.D. 2.1mm), (Ob2) objective for collecting scattering signals (100x, N.A. 0.90, W.D. 1.0mm), (NF) notch filter, (L1, L2, L3, LaS) convergent plano-convex achromatic lenses, (FM) flip mirror, (BS) 50/50 beamsplitter, ($_{aS_{1}}, BP$$_{aS_{2}}$) interference bandpass filter to anti-Stokes, (LS) aspheric condenser convergent lens, (Ob3) objective for laser focusing on APDS (20x, N.A. 0.45, W.D. 8.2-6.9 mm).
Figure 1 schematically describes the experimental setup used. A Verdi G-10 laser at 532 nm wavelength excites a Ti:Sapphire crystal from a Mira 900F laser, which generates a train of pulses at 76 MHz repetition rate, with interval between pulses of 13.6 ns and 130 fs nominal pulse width. The output signal, tuned at 781 nm wavelength with $104$ mW of power, passes through a half-wave plate (HWP) and a polarization beam splitter (PBS) to control the power and polarization before being focused on the diamond sample. The transmitted Raman signal is collected and a flippable mirror can direct this signal for two distinct optical paths. With the mirror flipped on the path, the signal is sent to a spectrometer and the Raman spectrum of the sample is obtained with a Charge Coupled Device (CCD). Once the mirror is flipped out of the optical path, the sample signal is divided in a 50/50 beamsplitter and each branch is separately sent through a HWP and a PBS, which are used to select the polarization of the photons. After polarization selection, in one of the branches a bandpass filter (FF01-650/150 - Semrock) and a longpass filter (LP02-633RE - Semrock) are used to filter the anti-Stokes photons that are then focused and detected in an avalanche photodiode (APD). In the other branch, the Stokes signal is filtered with a monochromator and then focused with an objective lens in an APD. The electric signals from the APDs (SPCM-AQRH-14 - Excelitas) are connected in a time-correlated single photon counting apparatus (PicoHarp 300 - PicoQuant) and a histogram of the temporal difference between the signals is created.
The sample used was a diamond grown by a CVD process (Type IIac, 100-oriented, from Almax) positioned so that the laser propagates in the (001) direction of the crystal. The sample assembly allows it to be rotated on the laser propagation axis such that the angle between the laser polarization and the crystallographic axes of the sample ($\theta$) can be varied. The polarization of the laser is vertical (V) with respect to the laboratory and the angle $\theta$ is set as $\theta=0^{\circ}$ for the (100) crystallographic axis of the sample matched with the vertical direction. For each monochromator position (each selected $\delta\omega=\omega_{c}-\omega_{S}$), one histogram of temporal difference is obtained in $300$ seconds of acquisition. This scanning procedure is done for two orientations of the sample ($\theta=0^{\circ}$ and $45^{\circ}$), and for each sample orientation the spectrum is obtained with the polarization of the photons incident on the avalanche photodiode (APD) detectors being selected in two ways: a spectrum with Stokes and anti-Stokes photons with vertical polarization (VV) and a spectrum with both photons with horizontal polarization (HH). Details of the treatment of the experimental data can be found in Appendixes A and B.
III.2 Results
@@
The intensity spectra of correlated SaS pairs for VV (blue dots) and HH (red dots) polarization at $\theta=0^{\circ}$ and $\theta=45^{\circ}$ are displayed in Figs. 2 (a) and (b), respectively. This experiment was specifically designed by combining the single-wavelength polarization analysis presented in [12] with the unpolarized spectral analysis provided in [15], to prove the strength of our theoretical approach. According to group theory analysis, VV(0) configuration does not exhibit a Raman contribution ($A^{R}_{xxxx}=0$ in Eq. (6)) and only involves non-resonant electronic transitions. For this reason, the blue curve shown in Fig. 2 (a) corresponding to this configuration is flat. Additionally, Fig. 2 (a) shows a dark blue solid line, obtained by averaging the intensity values $I_{SaS}^{corr}$ for this experimental configuration. We multiply this value by the monochromator slit spectral width $\Delta\omega/2\pi=11\text{ cm}^{-1}$, yielding $\langle I^{corr}_{SaS}(VV(0^{\circ})\rangle=(27.5\pm 3.5)\times 10^{3}\text{ Hz/W}^{2}$. Since $Y^{E}_{VV}(0^{\circ})=C\mathcal{E}_{0}^{2}A^{E}_{xxxx}$ and in our model it is proportional to $\langle I^{corr}_{SaS}(VV(0^{\circ}))\rangle^{1/2}$, we have a reference value for $A^{E}_{xxxx}$.
To compare the measured spectra with our spectral functions $f^{E}$ and $f^{R}$, Eq. (4), one must take into account the width of the frequency filters. The experimental setup (Fig. 1) contains a bandpass filter in the anti-Stokes branch, and a monochromator in the Stokes one. The anti-Stokes filter is modelled as a square frequency filter from anti-Stokes Raman shift of 900 cm-1 to 3000 cm-1, and the Stokes frequencies contributing to the counts are the ones that pass the monochromator slit, also modelled as a square frequency filter, but with width $\Delta\omega/2\pi=11\text{ cm}^{-1}$ around the Stokes central frequency $\omega_{S}$. We thus evaluate
$I^{\text{det}}_{li}(\omega_{S})=\int_{\omega_{S}-\Delta\omega/2}^{\omega_{S}+\Delta\omega/2}\left[\int_{F_{\text{aS}}}|\Psi_{li}(\omega_{S}^{\prime},\omega_{aS}^{\prime})|^{2}d\omega_{aS}^{\prime}\right]d\omega_{S}^{\prime},$ (7)
where $F_{\text{aS}}$ represents the aS filter frequencies, and $\Psi_{li}(\omega_{S},\omega_{aS})$ is the wave function in Eq. (II).
@@
By looking at the other relevant scattering geometries, we obtain the other tensorial components of Eq. (6) in relation to $A^{E}_{xxxx}$. The HH(45) configuration also exhibits only e-SaS, as seen in the red data in Fig 2 (b). The mean and standard deviation of the counts gives $(1.60\pm 0.28)\times 10^{3}$ Hz/W2. The HH(0) data in Fig. 2 (a) (red points) contain p-SaS with the characteristic resonance signature, which comes from $A^{R}_{xyyx}\neq 0$ in Eq. (6b), contributing with a Raman scattering in the $|HH\rangle$ polarization. Using $Y^{E}_{HH}(0^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(0.68-i0.12)$ and $Y^{R}_{HH}(0^{\circ})=51450\ [\text{Hz/W}^{2}]^{1/2}$ fits the data with the shown theoretical curve in Fig. 2 (a). The VV(45) data is the last configuration left, and it is shown in Fig. 2 (b) in blue. To fit it, we fix the Raman factor to be the same as in the HH(0) configuration, $Y^{R}_{VV}(45^{\circ})=51450\ [\text{Hz/W}^{2}]^{1/2}$, and use $Y^{E}_{VV}(45^{\circ})=\langle I_{SaS}^{corr}(\text{VV}(0^{\circ}))\rangle^{1/2}(1.61-i0.55)$. With the above values and working with Eq. (6) we obtain the electronic and Raman $A^{\eta}_{ijkl}$ components, summarized in Table 1. Details of the calculations can be found in Appendix C.
@@
Table 1: List of measured electronic and Raman tensor components ratios $A^{\eta}_{ijkl}/A^{E}_{xxxx}$.
@@
$(A^{E}_{xxyy}+A^{E}_{xyxy})/A^{E}_{xxxx}$ $(0.89\pm 0.10)+i(-0.07\pm 0.01)$
$A^{R}_{xxxx}/A^{E}_{xxxx}$ 0
@@
For completeness the second-order correlation function $g^{(2)}(0)$ is evaluated with the ratio $I^{corr}_{SaS}(\Delta\tau=0)+\bar{I}_{SaS}(\Delta\tau\neq 0)]/\bar{I}_{SaS}(\Delta\tau\neq 0)$, where $I^{corr}_{SaS}(\Delta\tau=0)$ are the measurements in Fig. 2 (a) and (b), and $\bar{I}_{SaS}(\Delta\tau\neq 0)$ accounts for the uncorrelated pair production. A plot of $g^{(2)}(0)$ is shown in Fig. 2(c,d), where it is evident that near the resonance the values are the lowest, accounting for the high number of uncorrelated SaS photon pairs produced. The large error bars in $g^{(2)}(0)$ reflect the errors in the low accidental counts (see Appendix A), and the theory agrees well with the data in spite of the error size. Furthermore, the curves for HH(0) and VV(45) show a drop in $g^{(2)}(0)$, being higher below and lower above the resonance region, while the VV($0^{\circ}$) and HH($45^{\circ}$) keep roughly the same value. The asymmetry in HH(0) and VV(45) is associated with the uncorrelated counts being symmetrical with respect to the resonance peak, and thus the asymmetry in the correlated counts for these configurations, which contain a p-SaS contribution, is transferred to the $g^{(2)}(0)$ curves. This asymmetry is a result of the constructive ($\delta\omega<\omega_{ph}$) versus destructive ($\delta\omega>\omega_{ph}$) interference between the e-SaS and the p-SaS, which explains the Cooper-pair-like behavior of $I_{SaS}^{corr}$ [10, 15]. On the other hand, the correlated counts in the VV(0) and HH(45) configurations are a purely e-SaS contribution, and thus are a flat curve, leading to an overall symmetric $g^{(2)}(0)$.
IV Polarization entanglement in the SaS pairs